What will be the velocity of the particle that is 1.8 m away from the centre on the minutes hand clock?
1
Expert's answer
2019-11-07T09:09:41-0500

The linear velocity


"v=\omega R=\frac{2\pi }{T}R"

"=\\frac{2\\pi }{3600\\:\\rm s}\\times 1.8\\:\\rm m=0.00314\u202c \\:m\/s=3.14\\: mm\/s"